OpenStudy (anonymous):

There are 18 bulls and 45 cows on a ranch. If 4 more bulls and 4 more cows were added, will the ratio of bulls to cow remain the same?

OpenStudy (anonymous):

Medal will be given :)

OpenStudy (anonymous):

no\[\frac{ 18 }{ 45 }\neq \frac{ 22 }{ 49 }\]

OpenStudy (anonymous):

Can you give me an explanation with words please it would be helpful @Euler271

OpenStudy (anonymous):

the ratio of A to B can be calculated \[\frac{ A }{ B }\] if this were true:\[\frac{ 18 }{ 45 } = \frac{ 18+4 }{ 45+4 }\] which isnt true

OpenStudy (anonymous):



OpenStudy (anonymous):

lol